OpenStudy (anonymous):

cos^2(x-pi/4)-3cosx=1

OpenStudy (anonymous):

@kirin what do we need to find here?

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

i think we need to solve \[\cos^2(x-\frac{ \pi }{ 4 })\] first we will use this formula cos(A-B) = cosA cosB +sinA sinB

OpenStudy (anonymous):

can you send answer ?

OpenStudy (anonymous):

sorry previous approach was wrong we are going to use half angle formulae cos2x = 2*cos^2x-1 it surely gonna give u an answer Sorry i cannot give direct answer here

OpenStudy (anonymous):

@kirin we can work step by step u just need to solve with me here

OpenStudy (dumbcow):

@niksva , does that approach give an exact answer? im not seeing it, this seems to have an numerical approximation solution

OpenStudy (anonymous):

This post is too difficult for me.I still do not solve it

OpenStudy (anonymous):

yeah it is not giving an exact answer we need to look at other way of solving this problem @dumbcow @kirin do u have any other idea @dumbcow ?

OpenStudy (dumbcow):

@kirin , is this calculus class?

OpenStudy (anonymous):

yes. this is my test

OpenStudy (dumbcow):

ok you can approximate using newtons method

OpenStudy (anonymous):

i will try

OpenStudy (dumbcow):

\[x_{n+1} = x_n - \frac{f(x_n)}{f'(x_n)}\] use \[f(x) = \sin(2x) -6\cos(x) -1\]

OpenStudy (anonymous):

@dumbcow r u suggesting newton raphson method here?

OpenStudy (dumbcow):

yes his approximation method for finding zeros

OpenStudy (anonymous):

this ques looks complicated

OpenStudy (anonymous):

The solutions to this are not nice at all

OpenStudy (anonymous):

yeahh it seems one has to play with numbers over here :)

OpenStudy (anonymous):

the solutions are irrational

OpenStudy (dumbcow):

there are 2 solutions within domain 0<x<2pi